OpenStudy (anonymous):

Find the rectangular equation for the conic. Identify it with focus, the center, the vertices, the directrix where it is applicable x=2+cosq y=3sinq-1

OpenStudy (dumbcow):

cosq = x-2 sinq = (y+1)/3 sin^2 + cos^2 = 1 --> (y+1)^2 /9 + (x-2)^2 = 1 looks like an ellipse